# Semantic-Consistent Bidirectional Contrastive Hashing for Noisy Multi-Label Cross-Modal Retrieval

### v1

@@ Introduction
The landscape of digital data has expanded from traditional text to multimodal content like images and videos, driven by internet advancements that complicate information retrieval. Cross-modal retrieval (Mandal and Biswas 2020; Hu et al. 2021) tackles this challenge by enabling searches across diverse modalities. For scalable and efficient large-scale retrieval across modalities, cross-modal hashing (Jiang and Li 2017) transforms heterogeneous data into a shared Hamming space through compact binary codes, allowing fast search with low computational overhead.
However, most existing methods rely on clean and fully annotated training data (Zhu et al. 2024). Such data are expensive to obtain in the real world, due to inherent data complexity (Frenay and Verleysen 2014) and human annotation errors (Song et al. 2022). While unsupervised methods reduce labeling costs, they often underperform compared to supervised approaches. A practical compromise is to use non-expert annotations, which inevitably introduce label noise, particularly in cross-modal multi-label scenarios.
@@
As illustrated in Figure 1, an image-text pair is annotated with three labels, one of which is noisy: the label “golf” is incorrect due to an annotation error. While robust learning strategies such as sample selection (Song et al. 2022) and loss correction (Wang et al. 2019) have been explored in single-modal tasks, they are not directly applicable to cross-modal settings due to modality heterogeneity. To address label noise in cross-modal retrieval while maintaining scalability and retrieval efficiency, several noise-robust hash learning methods (Wang et al. 2024; Shu et al. 2024; Pu et al. 2025) have been proposed. Although these methods aim to improve robustness, they often introduce new trade-offs. For example, (Shu et al. 2024) mitigates noise by excluding suspected noisy samples from training, which may inadvertently discard valuable supervisory signals. Meanwhile, (Wang et al. 2024) treats any pair with at least one shared label as a positive pair, and all others as negative. This binary treatment may overlook the nuanced semantic relationships common in multi-label scenarios, potentially resulting in imbalanced pair construction and underutilization of ambiguous yet informative samples.
@@
To address the challenge of learning with noisy labels, a variety of strategies have been proposed, which can be broadly categorized into two main approaches: robust learning and clean sample selection. For robust learning (Wang et al. 2025), since the conventional cross-entropy loss is sensitive to label noise, several studies have introduced noise-tolerant loss functions (Zhang and Sabuncu 2018) and regularization techniques (Liu et al. 2020; Wei et al. 2020) to enhance model robustness. For example, the symmetric cross-entropy loss (Wang et al. 2019) incorporates a noise-resistant objective that facilitates faster convergence and improved performance. However, empirical studies have shown that robust learning methods often face a trade-off between noise tolerance and predictive accuracy (Wang et al. 2022). In contrast, sample selection methods aim to identify clean instances from noisy datasets based on predefined criteria. For instance, MentorNet (Yu et al. 2019) introduces an auxiliary network to filter out noisy samples, while DivideMix (Li, Socher, and Hoi 2020) employs a Gaussian Mixture Model to dynamically divide training data into clean and noisy subsets. Nevertheless, these methods inevitably overlook hard examples with high loss that may still be informative, and they are less effective at bridging the modality gap in cross-modal tasks, as they are primarily designed for single-modality settings.
@@
Numerous cross-modal hashing methods have emerged, mainly categorized into unsupervised and supervised approaches. Unsupervised methods (Zhang et al. 2021; Zhu et al. 2022) learn a shared Hamming space by leveraging data structure without labels. For example, UCCH (Hu et al. 2023) incorporates contrastive learning with a triplet ranking loss to reduce modality gaps. However, the absence of supervision often limits their performance. In contrast, supervised methods (Wang et al. 2021; Wang and Peng 2022) assume that all training labels are clean and accurate, and generally optimize classification or similarity-preserving objectives, such as pairwise losses (Jiang and Li 2017) and triplet losses (Deng et al. 2018). This assumption, however, rarely holds in real-world multi-label scenarios where label noise is common. To tackle noisy supervision, robust methods like CMMQ (Yang et al. 2022) and NRCH (Wang et al. 2024) have been proposed. CMMQ introduces a proxy-based contrastive loss that prioritizes low-loss samples, while NRCH filters out unreliable samples using hard thresholds. However, such filtering risks discarding informative data, and most contrastive learning methods overlook semantic overlap in multi-label scenarios. These limitations highlight the necessity for more dynamic and noise-tolerant strategies for robust cross-modal hashing.
@@
To improve clarity, we first introduce the basic definitions and notations related to cross-modal hashing with noisy labels. Suppose the multimodal training dataset consists of $N$ examples, denoted as $\mathcal{D}=\left\{\left\{x_{i}^{m}\right\}_{m=1}^{2},\,y_{i}\right\}_{i=1}^{N}$, where $x_{i}^{m}$ represents the $m$-th modality of the $i$-th sample. Specifically, $m=1$ corresponds to the image modality and $m=2$ denotes the text modality. The label vector $y_{i}\in\mathbb{R}^{C}$ is potentially noisy, and $C$ denotes the total number of categories.
Cross-modal hashing learns modality-specific hash functions to embed heterogeneous data into a shared Hamming space. Let $f^{m}$ denote the hash function for modality $m$, where $m\in\{1,2\}$. During training, the continuous hash representation $h_{i}^{m}\in\mathbb{R}^{L}$ is computed as $h_{i}^{m}=\tanh(f^{m}(x_{i}^{m}))$, where $L$ is the length of the hash codes. The final binary hash codes $b_{i}^{m}\in\{-1,+1\}^{L}$ are obtained by applying the sign function, i.e., $b_{i}^{m}=\operatorname{sign}(h_{i}^{m})$.
In addition, a modality-specific linear classifier followed by a sigmoid activation, denoted as $g^{m}$, is used to produce a category-wise probability distribution. The predicted probability for sample $i$ in modality $m$ is computed as $z_{i}^{m}=g^{m}(h_{i}^{m})$, where $z_{i}^{m}\in[0,1]^{C}$.
@@
To fully utilize samples while learning from noisy sample interactions, we exploit cross-modal semantic consistency as a prior to assess label reliability: samples with semantically aligned cross-modal neighbors likely possess accurate labels. Based on this, we propose the cross-modal semantic-consistent classification (CSCC) module, which adaptively reweights samples by measuring their semantic proximity to neighbors. Firstly, we construct a soft label $p_{i}$ for each anchor sample by aggregating the labels of its most similar cross-modal neighbors:
$p_{i}=\sum_{k\in\mathcal{N}_{i}}y_{ik}\cdot(\frac{1}{2}\sum_{m=1}^{2}\frac{s_{ik}^{m}}{\sum_{j\in\mathcal{N}_{i}}s_{ij}^{m}})\quad,$ (1)
where $s^{m}_{ik}$ denotes the cosine similarity between the $i$-th sample in modality $m$ and its $k$-th neighbor, $y_{ik}$ represents the label vector of the $k$-th neighbor of sample $i$, and $\mathcal{N}_{i}$ denotes the set of neighbors of sample $i$.
@@
$w_{i}=\gamma+(1-\gamma)\cdot\frac{y_{i}\cdot p_{i}}{\|y_{i}\|\cdot\|p_{i}\|}\quad,$ (2)
@@
where $z_{i}^{m}$ denotes the predicted probability vector over $C$ classes for the $i$-th sample in modality $m$.
@@
Existing multi-label contrastive methods (Audibert, Gauffre, and Amini 2024; Zhang and Wu 2024) typically adopt either the ALL or ANY strategy (as shown in Figure 3) to construct positive and negative pairs: ALL requires identical labels for positives, while ANY considers any shared label as positive. All other pairs are treated as negatives. However, these binary schemes overlook the semantic nuances of partially overlapping labels in multi-label data.
To better reflect these nuanced relationships, we introduce a Bidirectional strategy (Figure 3) that categorizes sample pairs as positive (identical labels), negative (no shared labels), or soft (partial overlap), providing a finer-grained view of semantic similarity. To quantify label similarity using the Jaccard Index. Given two multi-hot label vectors ${y}_{i},{y}_{j}\in\{0,1\}^{C}$, the similarity score $R_{ij}$ is defined as:
$R_{ij}=\frac{{y}_{i}^{\top}{y}_{j}}{|{y}_{i}|_{1}+|{y}_{j}|_{1}-{y}_{i}^{\top}{y}_{j}}\quad,$ (4)
@@
where $S_{ij}=h^{1}_{i}\cdot h^{2}_{j}$ measures cross-modal feature similarity between samples $i$ and $j$, and $M_{ij}$ is a binary mask indicating positive pairs with $R_{ij}>0$, while $S_{ii}$ enforces intra-instance modality alignment as a regularizer. The hyperparameter $\xi$ controls the sharpness of positive similarity, guiding the model to perform fine-grained alignment based on label semantics for more coherent cross-modal fusion.
@@
To evaluate the performance of our method, we conducted extensive experiments across four widely-used benchmark datasets, NUS-WIDE (Chua et al. 2009), MS-COCO (Lin et al. 2014), MIRFlickr-25K (Huiskes and Lew 2008), and IAPR TC-12 (Escalante et al. 2010). And we conduct two cross-modal retrieval tasks: image-to-text (I2T) and text-to-image (T2I), where retrieval is based on Hamming distance. Similar to (Hu et al. 2023), we adopt the widely used Mean Average Precision (MAP) as the evaluation metric. MAP is computed as the mean of Average Precision (AP) scores over all queries, effectively reflecting both the precision and ranking quality of the retrieved results. MAP scores are reported under different hash code lengths (16, 32, 64, and 128 bits). To assess robustness, we inject mixed symmetric label noise (Burgert, Ravanbakhsh, and Demir 2022) into the training data at noise rates of 20%, 50%, and 80%.
@@
To validate the effectiveness of our proposed method, we conduct comprehensive comparisons against 12 state-of-the-art methods on four widely adopted cross-modal benchmark datasets. These baselines include five unsupervised approaches: DJSRH (Su, Zhong, and Zhang 2019), DGCPN (Yu et al. 2022), PIP (Zhang et al. 2021), CIRH (Zhu et al. 2022), and UCCH (Hu et al. 2023), as well as seven supervised methods: CPAH (Xie et al. 2020), CMMQ (Yang et al. 2022), MIAN (Zhang et al. 2023), LtCMH (Gao et al. 2023), NRCH (Wang et al. 2024), DHRL (Shu et al. 2024), and RSHNL (Pu et al. 2025). For each method, we track performance across training epochs and report the best MAP scores. Tables 1 and 2 report the average MAP scores for both I2T and T2I retrieval tasks. Additionally, Figure 4 plots the precision-recall curves under a 64-bit code length and 80% noise setting. Based on the results in Tables 1 and 2, we draw the following key observations:

### v2

@@ Introduction
The landscape of digital data has expanded from traditional text to multimodal content like images and videos, driven by internet advancements that complicate information retrieval. Cross-modal retrieval (Su et al. 2025b; Liu et al. 2024) tackles this challenge by enabling searches across diverse modalities. For scalable and efficient large-scale retrieval across modalities, cross-modal hashing (Jiang and Li 2017; Lan et al. 2025) transforms heterogeneous data into a shared Hamming space through compact binary codes, allowing fast search with low computational overhead.
However, most existing methods rely on clean and fully annotated training data (Zhu et al. 2024; Yin et al. 2025; Wu et al. 2025). Such data are expensive to obtain in the real world, due to inherent data complexity (Frenay and Verleysen 2014) and human annotation errors (Song et al. 2022). While unsupervised methods reduce labeling costs, they often underperform compared to supervised approaches. A practical compromise is to use non-expert annotations, which inevitably introduce label noise, particularly in cross-modal multi-label scenarios.
@@
As illustrated in Figure 1, an image-text pair is annotated with three labels, one of which is noisy: the label “golf” is incorrect due to an annotation error. While robust learning strategies such as sample selection (Song et al. 2022) and loss correction (Wang et al. 2019) have been explored in single-modal tasks, they are not directly applicable to cross-modal settings due to modality heterogeneity. To address label noise in cross-modal retrieval while maintaining scalability and retrieval efficiency, several noise-robust hash learning methods (Shu et al. 2024; Su et al. 2024; Cao et al. 2024) have been proposed. Although these methods aim to improve robustness, they often introduce new trade-offs. For example, (Shu et al. 2024) mitigates noise by excluding suspected noisy samples from training, which may inadvertently discard valuable supervisory signals. Meanwhile, (Wang et al. 2024) treats any pair with at least one shared label as a positive pair, and all others as negative. This binary treatment may overlook the nuanced semantic relationships common in multi-label scenarios, potentially resulting in imbalanced pair construction and underutilization of ambiguous yet informative samples.
@@
To address the challenge of learning with noisy labels, a variety of strategies have been proposed, which can be broadly categorized into two main approaches: robust learning and clean sample selection. For robust learning (Wang et al. 2025), since the conventional cross-entropy loss is sensitive to label noise, several studies have introduced noise-tolerant loss functions (Zhang and Sabuncu 2018) and regularization techniques (Liu et al. 2020; Wei et al. 2020) to enhance model robustness. For example, the symmetric cross-entropy loss (Wang et al. 2019) incorporates a noise-resistant objective that facilitates faster convergence and improved performance. However, empirical studies have shown that robust learning methods often face a trade-off between noise tolerance and predictive accuracy (Wang et al. 2022). In contrast, sample selection methods aim to identify clean instances from noisy datasets based on predefined criteria. For instance, MentorNet (Yu et al. 2019) introduces an auxiliary network to filter out noisy samples, while DivideMix (Li et al. 2020) employs a Gaussian Mixture Model to dynamically divide training data into clean and noisy subsets. Nevertheless, these methods inevitably overlook hard examples with high loss that may still be informative, and they are less effective at bridging the modality gap in cross-modal tasks, as they are primarily designed for single-modality settings.
@@
Numerous cross-modal hashing methods have emerged, mainly categorized into unsupervised and supervised approaches. Unsupervised methods (Zhang et al. 2021; Zhu et al. 2022) learn a shared Hamming space by leveraging data structure without labels. For example, UCCH (Hu et al. 2023) incorporates contrastive learning with a triplet ranking loss to reduce modality gaps. However, the absence of supervision often limits their performance. In contrast, supervised methods (Wang et al. 2021; Wang and Peng 2022; Su et al. 2025a) assume that all training labels are clean and accurate, and generally optimize classification or similarity-preserving objectives, such as pairwise losses (Jiang and Li 2017) and triplet losses (Deng et al. 2018). This assumption, however, rarely holds in real-world multi-label scenarios where label noise is common. To tackle noisy supervision, robust methods like CMMQ (Yang et al. 2022) and NRCH (Wang et al. 2024) have been proposed. CMMQ introduces a proxy-based contrastive loss that prioritizes low-loss samples, while NRCH filters out unreliable samples using hard thresholds. However, such filtering risks discarding informative data, and most contrastive learning methods overlook semantic overlap in multi-label scenarios. These limitations highlight the necessity for more dynamic and noise-tolerant strategies for robust cross-modal hashing.
@@
To improve clarity, we first introduce the basic definitions and notations related to cross-modal hashing with noisy labels. Suppose the multimodal training dataset consists of $N$ examples, denoted as $\mathcal{D}=\left\{\left\{\mathbf{x}_{i}^{m}\right\}_{m=1}^{2},\,\mathbf{y}_{i}\right\}_{i=1}^{N}$, where $\mathbf{x}_{i}^{m}$ represents the $m$-th modality of the $i$-th sample. Specifically, $m=1$ corresponds to the image modality and $m=2$ denotes the text modality. The label vector $\mathbf{y}_{i}\in\mathbb{R}^{C}$ is potentially noisy, and $C$ denotes the total number of categories.
Cross-modal hashing learns modality-specific hash functions to embed heterogeneous data into a shared Hamming space. Let $f^{m}$ denote the hash function for modality $m$, where $m\in\{1,2\}$. During training, the continuous hash representation $\mathbf{h}_{i}^{m}\in\mathbb{R}^{L}$ is computed as $\mathbf{h}_{i}^{m}=\tanh(f^{m}(\mathbf{x}_{i}^{m}))$, where $L$ is the length of the hash codes. The final binary hash codes $\mathbf{b}_{i}^{m}\in\{-1,+1\}^{L}$ are obtained by applying the sign function, i.e., $\mathbf{b}_{i}^{m}=\operatorname{sign}(\mathbf{h}_{i}^{m})$.
In addition, a modality-specific linear classifier followed by a sigmoid activation, denoted as $g^{m}$, is used to produce a category-wise probability distribution. The predicted probability for sample $i$ in modality $m$ is computed as $\mathbf{z}_{i}^{m}=g^{m}(\mathbf{h}_{i}^{m})$, where $\mathbf{z}_{i}^{m}\in[0,1]^{C}$.
@@
To fully utilize samples while learning from noisy sample interactions, we exploit cross-modal semantic consistency as a prior to assess label reliability: samples with semantically aligned cross-modal neighbors likely possess accurate labels. Based on this, we propose the cross-modal semantic-consistent classification (CSCC) module, which adaptively reweights samples by measuring their semantic proximity to neighbors. Firstly, we construct a soft label $\mathbf{p}_{i}$ for each anchor sample by aggregating the labels of its most similar cross-modal neighbors:
$\textbf{p}_{i}=\sum_{k\in\mathcal{N}_{i}}\mathbf{y}_{k}\cdot(\frac{1}{2}\sum_{m=1}^{2}\frac{s_{ik}^{m}}{\sum_{j\in\mathcal{N}_{i}}s_{ij}^{m}})\quad,$ (1)
where $s^{m}_{ik}$ denotes the cosine similarity between the $i$-th sample in modality $m$ and its $k$-th neighbor, $\mathbf{y}_{k}$ represents the label vector of the $k$-th neighbor of sample $i$, and $\mathcal{N}_{i}$ denotes the set of neighbors of sample $i$.
@@
$w_{i}=\gamma+(1-\gamma)\cdot\frac{\mathbf{y}_{i}\cdot\mathbf{p}_{i}}{\|\mathbf{y}_{i}\|\cdot\|\mathbf{p}_{i}\|}\quad,$ (2)
@@
where $\mathbf{z}_{i}^{m}$ denotes the predicted probability vector over $C$ classes for the $i$-th sample in modality $m$.
@@
Existing multi-label contrastive methods (Audibert et al. 2024; Zhang and Wu 2024) typically adopt either the ALL or ANY strategy (as shown in Figure 3) to construct positive and negative pairs: ALL requires identical labels for positives, while ANY considers any shared label as positive. All other pairs are treated as negatives. However, these binary schemes overlook the semantic nuances of partially overlapping labels in multi-label data.
To better reflect these nuanced relationships, we introduce a Bidirectional strategy (Figure 3) that categorizes sample pairs as positive (identical labels), negative (no shared labels), or soft (partial overlap), providing a finer-grained view of semantic similarity. To quantify label similarity using the Jaccard Index. Given two multi-hot label vectors $\mathbf{y}_{i},\mathbf{y}_{j}\in\{0,1\}^{C}$, the similarity score $R_{ij}$ is defined as:
$R_{ij}=\frac{\mathbf{{y}}_{i}^{\top}\mathbf{{y}}_{j}}{|\mathbf{{y}}_{i}|_{1}+|\mathbf{{y}}_{j}|_{1}-\mathbf{{y}}_{i}^{\top}\mathbf{{y}}_{j}}\quad,$ (4)
@@
where $S_{ij}=\mathbf{h}^{1}_{i}\cdot\mathbf{h}^{2}_{j}$ measures cross-modal feature similarity between samples $i$ and $j$, and $M_{ij}$ is a binary mask indicating positive pairs with $R_{ij}>0$, while $S_{ii}$ enforces intra-instance modality alignment as a regularizer. The hyperparameter $\xi$ controls the sharpness of positive similarity, guiding the model to perform fine-grained alignment based on label semantics for more coherent cross-modal fusion.
@@
To evaluate the performance of our method, we conducted extensive experiments across four widely-used benchmark datasets, NUS-WIDE (Chua et al. 2009), MS-COCO (Lin et al. 2014), MIRFlickr-25K (Huiskes and Lew 2008), and IAPR TC-12 (Escalante et al. 2010). And we conduct two cross-modal retrieval tasks: image-to-text (I2T) and text-to-image (T2I), where retrieval is based on Hamming distance. Similar to (Hu et al. 2023), we adopt the widely used Mean Average Precision (MAP) as the evaluation metric. MAP is computed as the mean of Average Precision (AP) scores over all queries, effectively reflecting both the precision and ranking quality of the retrieved results. MAP scores are reported under different hash code lengths (16, 32, 64, and 128 bits). To assess robustness, we inject mixed symmetric label noise (Burgert et al. 2022) into the training data at noise rates of 20%, 50%, and 80%.
@@
To validate the effectiveness of our proposed method, we conduct comprehensive comparisons against 12 state-of-the-art methods on four widely adopted cross-modal benchmark datasets. These baselines include five unsupervised approaches: DJSRH (Su et al. 2019), DGCPN (Yu et al. 2022), PIP (Zhang et al. 2021), CIRH (Zhu et al. 2022), and UCCH (Hu et al. 2023), as well as seven supervised methods: CPAH (Xie et al. 2020), CMMQ (Yang et al. 2022), MIAN (Zhang et al. 2023), LtCMH (Gao et al. 2023), NRCH (Wang et al. 2024), DHRL (Shu et al. 2024), and RSHNL (Pu et al. 2025). For each method, we track performance across training epochs and report the best MAP scores. Tables 1 and 2 report the average MAP scores for both I2T and T2I retrieval tasks. Additionally, Figure 4 plots the precision-recall curves under a 64-bit code length and 80% noise setting. Based on the results in Tables 1 and 2, we draw the following key observations:
@@ Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China (62306197, 62372315), Sichuan Science and Technology Planning Project (2025ZNSFSC1507, 2024YFG0007, 2024ZDZX0004, 2024NSFTD0049), Central Government’s Guide to Local Science and Technology Development Fund (2025ZYDF101), China Postdoctoral Science Foundation (2021TQ0223, 2022M712236), Chengdu Science and Technology Project (2023-XT00-00004-GX), Postdoctoral Joint Training Program of Sichuan University (SCDXLHPY2307).